OpenStudy (anonymous):

how do you simplify (b^4/5)(b^4/5) divided by b^3/5?

OpenStudy (anonymous):

(b^4/5)(b^4/5)=(b^4/5)^2=b^16/25 so your new expression would be (b^16/25)/(b^3/5) =(b^16/25)/(b^18/25) =1/(b^2/5)

OpenStudy (gw2011):

(b^4/5)(b^4/5)/b^3/5 = b^8/5/b^3/5 = b^5/5 = b